The probability of success and failure is 1. What could be the mathematics statistics interpretation of this
1
Expert's answer
2021-10-05T23:11:51-0400

We know that the sum of probability of success and failure is always equal to 1. That is,


"success+failure=1"

If the probability of success is 1 that means "P(S)=1" then,

probability of failure "P(F)= 1-1=0"


Vice-versa if the probability of success is 0 that means "P(S)=0" then,

probability of failure "P(F)=1-0=1"